# Language-Specific Latent Process Hinders Cross-Lingual Performance

### v2

@@ Abstract
Large language models (LLMs) are demonstrably capable of cross-lingual transfer, but can produce inconsistent output when prompted with the same queries written in different languages. To understand how language models are able to generalize knowledge from one language to the others, we apply the logit lens to interpret the implicit steps taken by LLMs to solve multilingual multi-choice reasoning questions. We find LLMs predict inconsistently and are less accurate because they rely on subspaces of individual languages, rather than working in a shared semantic space. While larger models are more multilingual, we show their hidden states are more likely to dissociate from the shared representation compared to smaller models, but are nevertheless more capable of retrieving knowledge embedded across different languages. Finally, we demonstrate that knowledge sharing can be modulated by steering the models’ latent processing towards the shared semantic space. We find reinforcing utilization of the shared space improves the models’ multilingual reasoning performance, as a result of more knowledge transfer from, and better output consistency with English.
@@ 1 Introduction
Humans have an innate ability to apply common knowledge and perform reasoning skills consistently across different languages, as the cognitive functions involved are largely independent of the languages they are learned in [5, 4]. Modern large language models (LLMs) are trained predominantly in English [24, 8] and Chinese [35]. To some degree, LLMs demonstrate cross-lingual capabilities [27], but the performance gap between languages remains [15, 13], and the models struggle to maintain consistent responses when they are probed by the same queries in different languages [20, 12, 7].
One possible explanation is that language models tend to develop language-specific neurons, rather than sharing them among multiple languages [29]. Nevertheless, language-specific neurons have never been shown to cause inconsistency in LLMs directly. Other work suggests that LLMs operate in a shared subspace closest to the English tokens, before projecting back to other languages in higher layers, for predictions in multilingual and multi-modal tasks [32, 34, 41]. These findings, however, fail to explain the lack of consistency across languages for the largest models [7] — otherwise, model outputs would converge toward English-prompted responses as multilingual performance improves.
@@
We first evaluate LLMs’ cross-lingual consistency and transfer on multi-choice queries. We find larger LLMs are indeed more accurate, but present large performance gaps, and remain inconsistent with poor cross-lingual transfer. We then investigate the intermediate steps taken implicitly by LLMs in solving the queries, and discover that models produce more divergent responses when operating in subspaces of different languages, rather than a shared internal representation that is closest to English hidden representation when processing multilingual prompts. We present evidence demonstrating a strong link between this phenomenon and the transfer performance of LLMs.
Implicit processing in English is more stable and accurate.
We interpret the hidden representations of LLMs using the logit lens [18, 32, 42, 34], which project intermediate hidden layers to the vocabulary space, to compute the latent distribution of answer choices. LLMs are capable of distinguishing multilingual answer choices in their hidden states, as indicated by our observation that the internal representations are closer to the embedding of the correct answers, relative to the other choices. Our results show the implicit process in English is more stable and more accurate than the process in the native languages of the prompts.
@@
In line with previous work, we find models of smaller scale work predominantly in the English shared space. Larger models, however, dissociate their internal states from the dominant semantic space, and move towards the individual subspaces of the native languages. While we have shown this phenomenon to undermine multilingual reasoning performance, the impact is less pronounced on larger models. We attribute this to their capability of encoding and retrieving information in different languages independently with significantly more parameters.
Inducing processing in the shared semantic space improves cross-lingual transfer and model performance in reasoning tasks.
Building on current understanding of knowledge transfer, how can we improve cross-lingual consistency while retaining performance of individual languages? The final part of this work demonstrates that knowledge sharing and consistency can be modulated by reinforcing or diminishing processing in the English latent space. To this end, we introduce cross-lingual activation steering, inspired by prior work [30, 23]. We find adding a steering vector towards English produces more English-like predictions, leading to performance gains in smaller models; whereas adding an opposite vector results in more divergent output, with a significant dip in performance.
@@ 2.1 Cross-lingual consistency and knowledge transfer
We define an LLM to be cross-lingually consistent when it produces the same output to the same queries written in different languages. Unlike Ifergan et al. [12], our measure of consistency is based upon multiple-choice question-answering tasks. To reduce language variability in the output space, we formulate the consistency between a language pair by their similarity in answer ranks, given by an LLM when prompted with the same questions and answer choices in their respective languages. Formally, for each of $i\in I$ questions in language $l$, $q_{i,l}\in Q_{l}$, paired with $j\in\{1..J\}$ answer choices $[a_{i,l}^{1},...,a_{i,l}^{J}]$, we rank the answer choices by their output probability $p(a_{i,l}^{j}|q_{i,l})$ given the question. We then combine the ranks $k(a_{i,l}^{j})$ of each answer within the set of choices:
$\mathbf{k}_{l}$ $=\langle k(a_{1,l}^{1}),...,k(a_{1,l}^{J}),k(a_{2,l}^{1}),...,k(a_{I,l}^{J})\rangle.$
Thus, the pairwise consistency between languages $l_{1}$ and $l_{2}$ is defined as the Spearman’s rank correlation coefficient between the rank vectors:
@@
Not all transfer between languages is beneficial—a model can be cross-lingually consistent but make many incorrect predictions. To approximate the rate of correct answers in a language propagates to another, we adapt our approach from Ifergan et al. [12] and estimate the positive transfer from $l_{1}$ to $l_{2}$ by the ratio of shared accurate responses to all accurate responses in $l_{1}$:
@@
$\operatorname{\mathbb{E}}[\textrm{consistency}]$ $=\frac{1}{|N|}\sum_{l_{1},l_{2}\in N}\mathrm{consistency}({l_{1},l_{2}}),$ (4)
@@
2.2 Interpreting the latent processes of multilingual tasks
@@
We rely on the logit lens to interpret the latent representation of a transformer by applying its output softmax layer to the hidden states [18]. These intermediate output distributions have been found to provide insights and largely explain the final predictions of a model [32, 1, 34].
@@
Logit lens is primarily designed to probe the latent distribution of the next token. To estimate the probability of a sequence in the latent space, we follow Zhong et al. [42] to allow past tokens to be included in the input, and decode from the hidden states to compute the final probabilities at each layer. Let $[x_{t},...,x_{T}]$ be the token indices of a phrase $\mathbf{w}$ following a query $\mathbf{q}$ with token indices $[x_{0},...,x_{t-1}]$, $h^{t}_{\ell}$ be the hidden state of layer $\ell$ at timestep $t$, and $O$ denote the embedding matrix. The latent probability of $\mathbf{w}$ at layer $\ell$, $p_{\ell}(\mathbf{w}|\mathbf{q})$, is equivalent to
$\prod_{\tau\in[t..T]}\mathrm{softmax}(Oh^{t}_{\ell})_{x_{\tau}}$ (7)
@@
which we normalize by the length of $\mathbf{w}$: $T-t+1]{p_{\ell}(\mathbf{w}|\mathbf{q})}$.111 The approach is analogous to prior work that analyzes the latent representation of LLMs, where traces of reasoning [36] and abstract concepts [14] are found, providing insights into model predictions. The proposed method is also plausible because it is directly related to how an autoregressive model predicts the next token, where a textual answer can be sampled from the token distribution.
Based on these equations, we operationalize the relative similarity between the hidden states and embedding space in terms of the normalized probabilities of the answer choices, which are evaluated in their full textual forms given a query. We then rank the answer choices accordingly and compute the consistency and transfer between all language pairs, using the methods described in Section 2.1. For clarity, we use $\eta^{\ell}_{m}$ to denote ranked answer choices in language $m$ queried in language $\eta$, derived from intermediate layer $\ell$. For example, ${\textrm{jv}}^{2}_{\alpha}$ and ${\textrm{jv}}^{2}_{\textrm{en}}$ refer to the ranked output at the second layer in alphabetical labels, and English respectively, from a Javanese query, which is to be distinguished from ${\textrm{jv}}^{2}_{\textrm{jv}}$ where the ranked output is provided in Javanese.
Using the latent probabilities and the ranked output, we test if language models could reason in languages beyond English and extract information stored in different languages [32, 1, 41, 34, 12]. Based on the findings of Tang et al. [29] and Goldman et al. [7], we hypothesize that large-scale LLMs, with a greater number of language-specific neurons in the hidden layers, could induce higher activation and utilization of multilingual subspace rather than the shared semantic space. We also expect operating in individual language subspace rather than the shared subspace to be the cause of limited knowledge transfer and inconsistency across multilingual prompts.
@@ 2.3 Cross-lingual activation steering
To further test these hypotheses, we derive from model behavior steering literature [30, 23] and propose cross-lingual activation steering. Because the shared latent space is most similar to English [32], we expect pushing the latent processes towards English to induce processing in the shared space, allowing for knowledge sharing. On the other hand, pushing the process towards the original language is expected to cause more divergent predictions.
Following Rimsky et al. [23], we obtain the contrastive addition vector of multiple-choice questions for each languages, by computing the mean difference in activation as the model processes parallel queries in English and each of the languages, for each data set. The vector is extracted from the last token of the prompts, as the models prepare to generate an answer in the next token. Specifically, for a set of parallel queries in English and language $m$, $Q_{{\textrm{en},m}}$, the steering vector at layer $\ell$ from $m$ to English is computed as:
$v^{\ell}_{m\rightarrow\textrm{en}}=\frac{1}{|Q_{{\textrm{en},m}}|}\sum_{q_{\textrm{en}},q_{m}\in Q_{{\textrm{en},m}}}h_{\ell}(q_{\textrm{en}})-h_{\ell}(q_{m}),$ (8)
where $h_{\ell}(...)$ refers to the hidden activation at layer $\ell$ of the query’s last token. Like previous work [23], we use a scalar multiplier $\gamma$ to modulate the direction of change and the magnitude of the vectors. $\gamma>0$ induces latent process similar to being prompted in English, whereas $\gamma<0$ reinforces language-specific processing. At inference, the extracted addition vector is added to the hidden activation as $\gamma\cdot v^{\ell}_{m\rightarrow\textrm{en}}$ at the last token of the prompts.
@@ 3 Experimental setup
We experiment on Gemma 2 [24] and Qwen 2.5 [35] family of language models, which have been trained and instruction-tuned primarily in English. We use 2B, 9B and 27B variants of Gemma models and 3B, 7B and 32B of Qwen models. Note that the Qwen models have been tuned on multilingual instructions and are claimed to support 29 languages [22]. For all models, we compute the output distribution of the hidden states at an interval of four layers, ending at the final layer (e.g., layer 2, 6, …, 42; where 42 is the total number of layers in Gemma 2 9B).
@@
We examine the latent process of the models using zero-shot prompts with three parallel multilingual tasks: Belebele for machine reading comprehension in 122 languages, created by human translators [2]; mMMLU for world knowledge and problem solving questions that have been manually translated and are available in 15 languages [9, 19]; and mHellaSwag for commonsense natural language inference in 10 languages, created with automatic translations [17, 39].
For multiple-choice questions, we associate each answer choice with an alphabetic characters (e.g., ‘A’), and evaluate the models based on the output distribution of the alphabets listed in the prompts. All prompt templates used are reported in Appendix A. We then derive based on Equation 7 the latent probabilities of the answer choices in i) their full textual form and ii) their parallel translations in English. We then follow Bandarkar et al. [2] to compute the accuracy of the models in these tasks.
@@ 4 Results
4.1 Cross-lingual performance and the impact of language relations
Table 1: Mean accuracy ($\uparrow$), consistency ($\uparrow$), positive ($\uparrow$) and negative ($\downarrow$) transfer. Larger models are more accurate, but the predictions are less consistent, with limited transfer across languages.
@@ 4.1 Cross-lingual performance and the impact of language relations
Gemma 2 2B 47.1$\pm$13.8 .381 .623 .372 42.5$\pm$5.9 .540 .671 .509 44.8$\pm$8.3 .461 .656 .470
Gemma 2 9B 72.4$\pm$18.4 .427 .779 .273 59.0$\pm$8.2 .561 .771 .478 66.2$\pm$11.9 .558 .780 .418
Gemma 2 27B 75.1$\pm$18.2 .548 .803 .289 62.6$\pm$8.3 .639 .810 .521 69.9$\pm$12.3 .631 .807 .429
[…]
Qwen 2.5 7B 57.6$\pm$20.3 .383 .650 .278 54.9$\pm$10.6 .541 .724 .442 60.4$\pm$15.0 .474 .718 .381
Qwen 2.5 32B 70.3$\pm$20.9 .485 .757 .277 66.4$\pm$11.8 .567 .798 .443 70.0$\pm$14.2 .555 .796 .396
@@
We report the accuracy and cross-lingual transfer performance of all models based on the ranked alphabet choices determined by the output distribution, i.e., between $x^{\lambda}_{\alpha}$ and $y^{\lambda}_{\alpha}$, $\lambda$ is the final layer, $x$ and $y$ are a language pair. Table 4.1 shows the mean accuracy across languages and the consistency and transfer rates (Equations 4, 5 and 6) for individual language models and benchmarks. Despite being more accurate, larger models such as Gemma 2 27B and Qwen 2.5 32B have more varied performances. In terms of consistency and transfer across languages, larger models offer little gain compared to their scale and accuracy, especially relative to the smaller models.
Figure 1 shows the pairwise positive transfer for Gemma 2 9B on mMMLU. Like Ifergan et al. [12], we find languages with known relations, such as Indic languages: Hindi (hi) and Bengali (bn); and Romance languages: Portuguese (pt), Spanish (es) and Italian (it); to be more closely connected and transfer more knowledge with each other. Indonesian (id) has higher rates of transfer with European languages, relative to Chinese (zh), Korean (ko) and Japanese (ja); indicating the role of shared scripts in knowledge transfer. The models are also less capable of generalizing predictions to and infer knowledge from low-resource languages: Swahili (sw) and Yoruba (yo). Knowledge learned in other languages remains inaccessible in yo as the models improve with increasing capacity.

### v3

@@ Abstract
Large language models (LLMs) are demonstrably capable of cross-lingual transfer, but can produce inconsistent output when prompted with the same queries written in different languages. To understand how language models are able to generalize knowledge from one language to the others, we measure representation similarity between languages, and apply the logit lens to interpret the implicit steps taken by LLMs to solve multilingual multi-choice reasoning questions. Our analyses reveal LLMs predict inconsistently and are less accurate because they rely on representations that are dissimilar across languages, rather than working in a shared semantic space. While larger models are more multilingual, we show their hidden states are more likely to dissociate from the shared representation compared to smaller models, but are nevertheless more capable of retrieving knowledge embedded across different languages. Finally, we demonstrate that knowledge sharing in small models can be facilitated by steering their latent processing towards the shared semantic space. This improves the models’ multilingual reasoning performance, as a result of more knowledge transfer from, and better output consistency with English.
@@ 1 Introduction
Humans have an innate ability to apply common knowledge and perform reasoning skills consistently across different languages, as the cognitive functions involved are largely independent of the languages they are learned in (Fodor, 1983; Dehaene, 2011). Modern large language models (LLMs) are trained predominantly in English (Rivière et al., 2024; Grattafiori et al., 2024) and Chinese (Yang et al., 2024a). To some degree, LLMs demonstrate cross-lingual capabilities (Shi et al., 2023), but the performance gap between languages remains (Kassner et al., 2021; Jiang et al., 2020), and the models struggle to maintain consistent responses when they are probed by the same queries in different languages (Qi et al., 2023; Ifergan et al., 2024; Goldman et al., 2025).
One possible explanation is that language models tend to develop language-specific neurons, rather than sharing them among multiple languages (Tang et al., 2024). Nevertheless, language-specific neurons have never been shown to cause inconsistency in LLMs directly. Other work suggests that LLMs operate in a shared subspace closest to the English tokens, before projecting back to other languages in higher layers, for predictions in multilingual and multi-modal tasks (Wendler et al., 2024; Wu et al., 2024; Zhao et al., 2024b). [^1] Wang et al. (2025a), for example, attribute language inconsistency to the models’ failure to transition from English to the target language at the final layers. These findings, however, fail to explain the lack of consistency across languages for the largest models (Goldman et al., 2025) — otherwise, model outputs would converge toward English-prompted responses as multilingual performance improves.
@@
We first evaluate LLMs’ cross-lingual consistency and transfer on multi-choice queries. We find larger LLMs are indeed more accurate, but present large performance gaps, and remain inconsistent with subpar cross-lingual transfer. Through measuring cross-lingual alignment in representations, we discover that models produce more divergent responses when their internal representations are more dissimilar across languages. We present evidence demonstrating a strong link between utilization of a shared semantic space and the transfer performance of LLMs.
@@
Despite their superior multilingual performance, we find representations in larger LLMs to be more structurally dissimilar and farther apart in the semantic space across languages. Further analyses with logit lens reveal that the representations move towards individual subspaces of the native languages. In contrast, smaller models tend to work predominantly in the English shared space. We attribute this to larger models’ capability of encoding and retrieving information in different languages independently with significantly more parameters.
Inducing processing in the shared semantic space improves cross-lingual transfer and performance in smaller models.
To provide a causal evidence for our claims, we apply cross-lingual steering in the final part of this work, inspired by Turner et al. (2023); Rimsky et al. (2024). These linear steering vectors are constructed to reinforce latent processing in the shared space, leading to more English-like predictions and performance gains. We find adding a linear steering vector towards English is effective for small models, but not for larger ones, likely due to their increased non-linearity and structurally distinct language subspaces.
@@ 2.1 Cross-lingual consistency and knowledge transfer
We define an LLM to be cross-lingually consistent when it produces the same output to the same queries written in different languages. Unlike Ifergan et al. (2024), our measure of consistency is based upon multiple-choice question-answering (MCQ) tasks. To reduce language variability in the output space, we formulate the consistency between a language pair by their similarity in answer ranks, given by an LLM when prompted with the same questions and answer choices in their respective languages. MCQ ranks are informative, as the distractors are typically selected to align to feasible mistakes, which we expect to transfer across languages.
Formally, for each of $i\in I$ questions in language $l$, $q_{i,l}\in Q_{l}$, paired with $j\in\{1..J\}$ answer choices $a_{i,l}^{1},...,a_{i,l}^{J}]$, we rank the answer choices by their output probability $p(a_{i,l}^{j}|q_{i,l})$ given the question. We then combine the ranks $k(a_{i,l}^{j})$ of each answer within the set of choices:
[…]
Thus, the pairwise consistency between languages $l_{1}$ and $l_{2}$ is defined as the Spearman’s rank correlation coefficient between the rank vectors: [^2] We choose Spearman’s correlation based on ranks over Pearson’s correlation based on answer distributions because the former is less sensitive to variance introduced by multilingual textual outputs. We decide against KL divergence for similar reason, and because the concept of consistency is more intuitive than inconsistency.
@@
Not all transfer between languages is beneficial—a model can be cross-lingually consistent but make many incorrect predictions. To approximate the rate of correct answers in a language propagates to another, we adapt our approach from Ifergan et al. (2024) and estimate the positive transfer from $l_{1}$ to $l_{2}$ by the ratio of shared accurate responses to all accurate responses in $l_{1}$:
@@
[…]
2.2 Cross-lingual alignment and multilingual processing in the latent space
Like Hämmerl et al. (2024), we hold the view that cross-lingual alignment is a property where models represent queries in different languages with equivalent semantics more similarly than those with dissimilar semantics. We choose two metrics to quantify cross-lingual similarity in representations: linear CKA and cosine similarity, which we justify below. To further provide insights on the ways in which representations are dissimilar, we follow prior work and use logit lens, which estimate the similarity between the hidden representation and the language-specific embedding space (Wendler et al., 2024; Alabi et al., 2024).
Linear CKA.
CKA is proposed as a similarity measure between representations and is invariant to any orthogonal transformations and isotropic scaling (Kornblith et al., 2019). These properties make them ideal for determining the structural similarity between neural networks that have different random initializations and widths. We argue that CKA could also be useful for detecting any meaningful similarity between parallel representations of different languages and for comparison across models at different scales. Following Kornblith et al. (2019), we apply a linear kernel for all our CKA measures, unless otherwise specified.
Formally, let $X=[h^{l_{1}}_{1,\ell},...,h^{l_{1}}_{n,\ell}]$, $Y=[h^{l_{2}}_{1,\ell},...,h^{l_{2}}_{n,\ell}]$ be the representations of $n$ parallel queries in languages $l_{1}$ and $l_{2}$ for layer $\ell$. The linear CKA between languages is defined as
[…]
Cosine similarity.
Cosine similarity is a widely used metric for cross-lingual alignment (Hämmerl et al., 2024; Chua et al., 2024; Wang et al., 2025a). We compute the cosine similarity between a language pair by the mean similarity over a set of parallel queries:
[…]
The function captures the degree to which two representations belong in the same neighborhood, but is insensitive to the structural relations between examples (Linzen, 2016; Zhou et al., 2022), unlike CKA. Further, cosine similarity inherently favors high-dimensional space, which is undesirable for our purpose of comparing between smaller and larger models with different hidden dimensions. To mitigate this effect, we take inspiration from Hämmerl et al. (2024) and compute language-pair similarity as its ratio to the similarity between representations of different semantics in the monolingual space of either language. This can be obtained by repeated sampling from the same query set, provided that the paired samples are different to each other:
[…]
Finally, we define the normalized cosine similarity between languages as the harmonic mean of the normalizations by $\mathrm{cos}(l_{1},\ell)$ and $\mathrm{cos}(l_{2},\ell)$:
$\mathrm{cos}_{\mathrm{norm}}(l_{1},l_{2},\ell)=\frac{2\mathrm{cos}_{l_{1}}\mathrm{cos}_{l_{2}}}{\mathrm{cos}_{l_{1}}+\mathrm{cos}_{l_{2}}},$ (10)
$\textrm{where}\quad\mathrm{cos}_{l_{1}}=\mathrm{cos}(l_{1},l_{2},\ell)/\mathrm{cos}(l_{1},\ell)\quad\textrm{and}\quad\mathrm{cos}_{l_{2}}=\mathrm{cos}(l_{1},l_{2},\ell)/\mathrm{cos}(l_{2},\ell).$ (11)
Logit lens.
@@ 2.2 Cross-lingual alignment and multilingual processing in the latent space
To characterize the cross-lingual differences in representation, we rely on the logit lens to interpret the hidden states by applying the output softmax layer to the vectors (Nostalgebraist, 2020). These intermediate distributions have been found to provide insights and largely explain the final predictions of a model (Wendler et al., 2024; Alabi et al., 2024; Wu et al., 2024).
@@
Logit lens is primarily designed to probe the latent distribution of the next token. To estimate the probability of a sequence in the latent space, we follow Zhong et al. (2024) to allow past tokens to be included in the input, and decode from the hidden states to compute the final probabilities at each layer. Let $x_{t},...,x_{T}]$ be the token indices of a phrase $\mathbf{w}$ following a query $\mathbf{q}$ with token indices $x_{0},...,x_{t-1}]$, $h^{t}_{\ell}$ be the hidden state of layer $\ell$ at timestep $t$, and $U$ denote the unembedding matrix. The latent probability of $\mathbf{w}$ at layer $\ell$, $p_{\ell}(\mathbf{w}|\mathbf{q})$, is equivalent to
[…]
@@
which we normalize by the length of $\mathbf{w}$: $\sqrt[T-t+1]{p_{\ell}(\mathbf{w}|\mathbf{q})}$. [^3] The approach is analogous to prior work that analyzes the latent representation of LLMs, where traces of reasoning (Yang et al., 2024b) and abstract concepts (Jin et al., 2025) are found, providing insights into model predictions. The proposed method is also plausible because it is directly related to how an autoregressive model predicts the next token, where a textual answer can be sampled from the token distribution.
Based on these equations, we operationalize the relative similarity between the hidden states and embedding space in terms of the normalized probabilities of the answer choices in full textual forms. For all queries, we compute the probabilities of the answer choices in the native language, as well as that of English parallel choices. Using the latent probabilities, we test if language models could reason in languages beyond English and extract information stored in different languages (Wendler et al., 2024; Alabi et al., 2024; Zhao et al., 2024b; Wu et al., 2024; Ifergan et al., 2024).
Cross-lingual alignment is assumed to underlie zero-shot cross-lingual transfer for many applications (Hämmerl et al., 2024). Therefore, our first hypothesis is that languages with more similar representations to those of other languages will also be more consistent. Based on Tang et al. (2024) and Goldman et al. (2025), we also hypothesize that large-scale LLMs, with a greater number of language-specific neurons in the hidden layers, could induce more diverging representations across languages. Overall, we expect models that are more cross-lingually aligned to operate in the shared semantic space, which is closest to English, rather than individual language subspace, facilitating cross-lingual knowledge transfer and consistency.
@@ 2.3 Cross-lingual activation steering
To further test these hypotheses, we derive from model behavior steering literature (Turner et al., 2023; Rimsky et al., 2024) and propose cross-lingual activation steering. Because the shared latent space is most similar to English (Wendler et al., 2024), we expect pushing the latent processes towards English to induce processing in the shared space, allowing for knowledge sharing, whereas reversing the process will cause more divergent predictions.
Following Rimsky et al. (2024), we obtain the contrastive addition vector of multiple-choice questions for each languages, by computing the mean difference in activation as the model processes parallel queries in English and each of the languages, for each data set. The vector is extracted from the last token of the prompts, as the models prepare to generate an answer in the next token. Specifically, for a set of parallel queries in English and language $m$, $Q_{{\textrm{en},m}}$, the steering vector at layer $\ell$ from $m$ to English is computed as:
[…]
@@ 3 Experimental setup
We experiment on Gemma 2 (Rivière et al., 2024) and Qwen 2.5 (Yang et al., 2024a) family of language models, which have been trained and instruction-tuned primarily in English. We use 2B, 9B and 27B variants of Gemma models and 3B, 7B and 32B of Qwen models. Note that the Qwen models have been tuned on multilingual instructions and are claimed to support 29 languages (Qwen Team, 2024). For all models, we compute the output distribution of the hidden states at an interval of four layers, ending at the final layer (e.g., layer 2, 6, …, 42; where 42 is the total number of layers in Gemma 2 9B).
@@
We examine the latent process of the models using zero-shot prompts with three parallel multilingual tasks: Belebele for machine reading comprehension in 122 languages, created by human translators (Bandarkar et al., 2023); mMMLU for world knowledge and problem solving questions that have been manually translated and are available in 15 languages (Hendrycks et al., 2020; OpenAI, 2024); and mHellaSwag for commonsense natural language inference in 10 languages, created with automatic translations (Lai et al., 2023; Zellers et al., 2019).
For multiple-choice questions, we associate each answer choice with an alphabetic characters (e.g., ‘A’). The model output is constrained to these characters to ensure any observations made in the latent space are not due to variance in multilingual output. [^4] With our setup, it is possible for models to achieve perfect consistency with maximally aligned representations, and any variation we observe can solely be attributed to the differences in multilingual processing. We then evaluate the models based on the output distribution of the alphabets listed in the prompts. All prompt templates used are reported in Appendix A. We follow Bandarkar et al. (2023) to compute the accuracy of the models in these tasks.
@@ 4 Results
Table 1: Mean accuracy ($\uparrow\uparrow$), consistency ($\uparrow\uparrow$), positive ($\uparrow\uparrow$) and negative ($\downarrow\downarrow$) transfer. Larger models are more accurate, but with larger performance gap and limited transfer across languages.
@@
Gemma 2 2B 47.113.8 .381 .623 .372 42.55.9 .540 .671 .509 44.88.3 .461 .656 .470
Gemma 2 9B 72.418.4 .427 .779 .273 59.08.2 .561 .771 .478 66.211.9 .558 .780 .418
Gemma 2 27B 75.118.2 .548 .803 .289 62.68.3 .639 .810 .521 69.912.3 .631 .807 .429
Qwen 2.5 3B 51.717.9 .276 .599 .278 47.89.1 .423 .662 .440 52.914.1 .418 .663 .385
Qwen 2.5 7B 57.620.3 .383 .650 .278 54.910.6 .541 .724 .442 60.415.0 .474 .718 .381
Qwen 2.5 32B 70.320.9 .485 .757 .277 66.411.8 .567 .798 .443 70.014.2 .555 .796 .396
@@
We report the accuracy and cross-lingual transfer performance of all models based on the ranked alphabet choices. Table 1 shows the mean accuracy across languages and the consistency and transfer rates (Equations 4, 5 and 6) for individual language models and benchmarks. Despite being more accurate, larger models such as Gemma 2 27B and Qwen 2.5 32B have more varied performances. In terms of consistency and transfer across languages, larger models offer little gain compared to their scale and accuracy, especially relative to the smaller models. Like Ifergan et al. (2024), we also found the impact of language relations on cross-lingual transfer, which we discuss in Appendix C.
4.1 Cross-lingually similar representations predict knowledge transfer
(a) linear CKA
(b) cosine similarity (normalized)
Figure 1: Mean similarity between representations of language pairs. Latent representations of Gemma 9B and Qwen 7B are the most aligned cross-lingually relative to their smaller and larger counterparts (lighter and darker blues respectively), especially towards the middle layers. Shaded areas indicate 95% confidence intervals across language pairs. Note that each model consists of different number of layers, which we normalize by the maximum layer along x-axes.
Table 2: Pearson’s correlations between averaged cross-lingual similarity measures of languages and their output accuracy, consistency and positive transfer from other languages. tr+ refers to the mean positive transfer from all source languages (Equation 2). * indicates $p<.05$, ** indicates $p<.01$ and ***, $p<.001$.
Belebele mMMLU mHellaSwag
acc. cons. tr+ acc. cons. tr+ acc. cons. tr+
linear CKA
Gemma 2 2B 0.760.76∗∗∗ 0.850.85∗∗∗ 0.810.81∗∗∗ 0.640.64∗∗ 0.950.95∗∗∗ 0.810.81∗∗∗ $0.72$ 0.900.90∗∗∗ 0.880.88∗∗∗
Gemma 2 9B 0.940.94∗∗∗ 0.950.95∗∗∗ 0.940.94∗∗∗ 0.940.94∗∗∗ 0.990.99∗∗∗ 0.970.97∗∗∗ 0.930.93∗∗∗ 0.990.99∗∗∗ 0.970.97∗∗∗
Gemma 2 27B 0.970.97∗∗∗ 0.970.97∗∗∗ 0.970.97∗∗∗ 0.950.95∗∗∗ 0.980.98∗∗∗ 0.970.97∗∗∗ 0.900.90∗∗∗ 0.970.97∗∗∗ 0.950.95∗∗∗
Qwen 2.5 3B 0.440.44∗∗∗ 0.560.56∗∗∗ 0.450.45∗∗∗ 0.790.79∗∗∗ 0.860.86∗∗∗ 0.830.83∗∗∗ 0.840.84∗∗ 0.950.95∗∗∗ 0.890.89∗∗∗
Qwen 2.5 7B 0.610.61∗∗∗ 0.600.60∗∗∗ 0.630.63∗∗∗ 0.860.86∗∗∗ 0.890.89∗∗∗ 0.870.87∗∗∗ 0.920.92∗∗∗ 0.990.99∗∗∗ 0.960.96∗∗∗
Qwen 2.5 32B 0.740.74∗∗∗ 0.750.75∗∗∗ 0.750.75∗∗∗ 0.680.68∗∗ 0.710.71∗∗ 0.710.71∗∗ 0.940.94∗∗∗ 0.960.96∗∗∗ 0.960.96∗∗∗
cosine similarity (normalized)
Gemma 2 2B 0.680.68∗∗∗ 0.820.82∗∗∗ 0.740.74∗∗∗ 0.760.76∗∗∗ 0.960.96∗∗∗ 0.900.90∗∗∗ $0.69$ 0.880.88∗∗∗ 0.850.85∗∗
Gemma 2 9B 0.930.93∗∗∗ 0.960.96∗∗∗ 0.940.94∗∗∗ 0.900.90∗∗∗ 0.970.97∗∗∗ 0.940.94∗∗∗ 0.910.91∗∗∗ 0.980.98∗∗∗ 0.960.96∗∗∗
Gemma 2 27B 0.780.78∗∗∗ 0.810.81∗∗∗ 0.790.79∗∗∗ 0.870.87∗∗∗ 0.850.85∗∗∗ 0.870.87∗∗∗ 0.920.92∗∗∗ 0.970.97∗∗∗ 0.960.96∗∗∗
Qwen 2.5 3B 0.330.33∗∗∗ 0.340.34∗∗∗ 0.340.34∗∗∗ $0.47$ 0.660.66∗∗ $0.56$ $0.51$ $0.65$ $0.57$
Qwen 2.5 7B 0.560.56∗∗∗ 0.650.65∗∗∗ 0.580.58∗∗∗ 0.840.84∗∗∗ 0.890.89∗∗∗ 0.860.86∗∗∗ 0.820.82∗∗ 0.930.93∗∗∗ 0.880.88∗∗∗
Qwen 2.5 32B 0.330.33∗∗∗ 0.360.36∗∗∗ 0.340.34∗∗∗ $0.63$ $0.62$ 0.650.65∗∗ $0.73$ 0.810.81∗∗ 0.780.78∗∗
We compute cross-lingual alignment by CKA and cosine similarity (Equations 7 and 10), based on 50 randomly sampled, parallel queries from each data set. Figure 1 shows the similarity values across layers for all models. CKA reaches the peak values in the lower layers then decreases. In contrast, cosine similarity in the early layers suggests little alignment, but continues to increase at higher layers (< 0.8). This indicates multilingual inputs are initially encoded in a structurally similar way, but are far apart in latent space. In deeper layers, however, the representations move towards the same direction, but become less similar in terms of geometric structure. Nevertheless, we observe in both cases that middle-sized models (Gemma 9B and Qwen 7B) are more cross-lingually aligned than the smaller and larger models. This partially supports our hypothesis, and occurs in spite of the largest models (Gemma 27B and Qwen 32B) attaining the best performance in multilingual tasks. It is thus possible that the models could encode and retrieve more accurately knowledge learned in different languages independently with higher capacity. Representations then diverge closest to the output layer, as the models prepare to generate the next token in the native language.
Do models predict more accurately and consistently with shared representation? Table 2 reports the Pearson’s correlations between the similarity measures (averaged across layers) extracted from 50 examples, and accuracy, consistency and positive transfer across languages of held-out examples. Overall, the results demonstrate strong predictive power of the similarity meausres on transfer performance, and that languages with more similar representations to the others are more likely to receive positive transfer, make more consistent predictions, and are therefore, more accurate.